4. Investors expect Microtech to pay the first dividend of $1.25 at the end of Year 2. The dividend should grow at a rate of 8 percent per year during Years 3, 4, and 5. The stock can be sold for $42 at the end of Year 5. Determine the current market value of the company’s stock if the required rate of return is 12 percent.

$27.61 Please show all work typed as I am on mobile
Answer #1

Returns from stock = returns from dividend + returns from capital gains

Lets map the cash flows first;All units in dollar expect %s

Year 1=0

Year 2=1.25

Year 3= 1.08*1.25=1.35

Year 4=1.35*1.08=1.458

Year 5=1.458*1.08=1.574 + 42 (on sale of share)=43.574

Reqd rate of return = 12%

So we have to discount the cashflows accordingly

[0/(1.12)]+[1.25/(1.12^2)]+[1.35/(1.12^3)]+[1.458/(1.12^4)]+[43.574/(1.12^5)]=$27.61